OpenStudy (anonymous):

Write the equation of a line in slope intercept form that passes through the points (0, 7) and (5, 3)?

OpenStudy (anonymous):

First find the slope

OpenStudy (anonymous):

sorry, i no see you. ): so i find slope? how?

OpenStudy (anonymous):

i no no how to see slope..

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) thn put in the eq. y=mx+c

OpenStudy (anonymous):

i understan little. tank you help me! :)